Black iron oxide is an ore containing iron and oxygen that occurs in magnetite. A 2.4480 g sample of the ore is found to contain 1.7714g of iron. Calculate the empirical formula of black iron oxide.​

1
Expert's answer
2022-02-10T22:54:03-0500

Mass of ore= 2.4480g

Iron= 1.7714

Mass of oxygen= 0.6766

Moles= mass /rmm

Moles of Fe= 1.7714/55.8= 0.032

Moles of O= 0.6766/16= 0.042

Removing decimal by a common factor × 100

Fe= 3

O= 4

Formula= Fe3O4